OpenStudy (anonymous):

Is the sequence 5, 9, 15, an arithmetic sequence? Explain

jimthompson5910 (jim_thompson5910):

if you can show that the difference between successive terms is the same, then it's an arithmetic sequence

OpenStudy (anonymous):

Subtract T1 from T2 and T2 From T3 if the difference is the same then it is .

jimthompson5910 (jim_thompson5910):

ex: 1, 5, 9, 13, ... is arithmetic since (5-1=4, 9-5 = 4, 13-9 = 4...all differences are 4) but 1, 7, 12, 14 is NOT arithmetic (7-1 = 6, 12 - 7 = 5, 14 - 12 = 2...the differences aren't the same)

OpenStudy (precal):

arithmetic means that you can add a number or subtract a number throughout the sequence (just like jim_thompson5910 stated above)